OpenStudy (anonymous):

2x+4=-3x-11 Now which step do I do first to isolate the variable?

OpenStudy (anonymous):

Move -3x to LHS and 4 to RHS

OpenStudy (anonymous):

add 3x to both sides to give 5x + 4 = 11 then subtract 4 from both sides 5x = -15

OpenStudy (angela210793):

2x+3x=-11-4 (+3x and -4 both sides) 5x=-15 (divide by 5) x=15/5 x=-3

OpenStudy (anonymous):

1.Subtract 3x from both sides 2.Add 3x to both sides 3.Add 2x to both sides

OpenStudy (anonymous):

2? Right

OpenStudy (anonymous):

2.